### v1

# Duistarmaat’s triple index and the difference of Hermitian matrices

Abstract.
In this paper we develop a systematic calculus for the Duistermaat index, a symplectic invariant defined for triples of Lagrangian subspaces. Introduced nearly half a century ago, this index has lately been the subject of renewed attention, due to its central role in eigenvalue interlacing problems on quantum graphs (and more abstractly for self-adjoint extensions of symmetric operators). Here we give an axiomatic characterization of the index that leads to elementary proofs of its fundamental properties. We also relate the index to other quantities often appearing in symplectic geometry, such as the Hörmander–Kashiwara–Wall index and the Maslov index. Among other things, this leads to a curious formula for the Morse index of a difference of Hermitian matrices.
1. Introduction
An important problem in the spectral theory of metric graphs is to quantify changes in the Laplacian eigenvalues when a finite number of vertex conditions are altered. [^1] For a survey of results when the vertex conditions are of a particularly well-studied type, we refer the reader to [5, Sec. 3.1]. Abstractly, this amounts to comparing self-adjoint extensions of a symmetric operator with finite and equal defect numbers. It was recently observed [4] that the difference of eigenvalue counting functions can be efficiently bounded in terms of a certain topological index.
The relevant index, introduced by Duistermaat in [9], takes three Lagrangian subspaces as its input. In the metric graph setting these correspond to the two sets of vertex conditions under consideration and also the Dirichlet conditions at the affected vertices. Duistarmaat’s definition is somewhat indirect, as it involves an arbitrary choice of a fourth Lagrangian plane that is transversal to the three planes of interest. The proof that the resulting index is independent of this choice relies on the Maslov index for paths of Lagrangian subspaces, see [9, Eq. (2.16)]. The Maslov index is an invariant depending on the homotopy class of the path (not just its endpoints), and hence it is inherently more complicated than an index depending on fixed Lagrangian subspaces.
The principal aim of this note is to define the Duistermaat index directly, without recourse to the Maslov index. We do this by giving a set of axioms (inspired by the approach of [8]) as well as a formula requiring nothing more than linear algebra. It is the set of axioms that turns out to be most useful in computation, allowing us to give short and simple proofs for many known and some unknown properties of the Duistermaat index.
One application of the developed theory is the following curious identity and its generalizations: for any invertible $n\times n$ Hermitian matrices $A$ and $B$,
$n_{-}\big{(}A-B\big{)}-n_{-}\big{(}B^{-1}-A^{-1}\big{)}=n_{-}(A)-n_{-}(B),$ (1.1)
where $n_{-}(X)$ denotes the number of negative eigenvalues (also known as the Morse index) of the matrix $X$. It is natural to wonder what happens if either $A$ or $B$ fails to be invertible. Some special cases are described Proposition 5.1, when $\ker A\subseteq\ker B$ or vice versa. The following version, however, is valid for any $n\times n$ Hermitian matrices $A$ and $B$:
$n_{-}\big{(}A-B\big{)}=n_{-}(A)-n_{-}(B)+\operatorname{\iota}({\mathcal{G}_{0}},\mathcal{G}_{B},\mathcal{G}_{A}),$ (1.2)
where $\mathcal{G}_{M}$ stands for the graph of the matrix $M$,
$\mathcal{G}_{M}:=\big{\{}(x,Mx):x\in\mathbb{C}^{n}\big{\}},$ (1.3)
and $\operatorname{\iota}$ is the Duistermaat index, which we are about to define.
We consider $\mathbb{C}^{n}\oplus\mathbb{C}^{n}$ as a Hermitian symplectic space with the canonical symplectic form
$\omega\big{(}(x_{1},x_{2}),(y_{1},y_{2})\big{)}:=\left<x_{1},y_{2}\right>_{\mathbb{C}^{n}}-\left<x_{2},y_{1}\right>_{\mathbb{C}^{n}}.$ (1.4)
A subspace $\mathcal{L}\subset\mathbb{C}^{n}\oplus\mathbb{C}^{n}$ is Lagrangian if it coincides with its symplectic complement,
$\mathcal{L}^{\omega}:=\left\{v\in\mathbb{C}^{n}\oplus\mathbb{C}^{n}\colon\omega(u,v)=0\text{ for all }u\in\mathcal{L}\right\}.$ (1.5)
The set of Lagrangian subspaces (often called Lagrangian planes) in $(\mathbb{C}^{n}\oplus\mathbb{C}^{n},\omega)$ is denoted by $\Lambda(n)$. Every Lagrangian subspace has dimension $n$, half of the dimension of the symplectic space, therefore $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ are transversal whenever $\mathcal{L}_{1}\cap\mathcal{L}_{2}=\{0\}$.
An important example of a Lagrangian plane is the graph of a Hermitian matrix, as in (1.3). In particular, the “horizontal” plane,
${\mathcal{G}_{0}}:=\mathbb{C}^{n}\oplus 0,$ (1.6)
will play an important role in this note. The “vertical” plane $0\oplus\mathbb{C}^{n}$ is not the graph of a Hermitian matrix, even though we will abuse the notation in (1.3) to denote it
${\mathcal{G}_{\infty}}:=0\oplus\mathbb{C}^{n}.$ (1.7)
Being “somewhat vertical” is the only obstacle for a Lagrangian plane to be a graph: a Lagrangian plane $\mathcal{L}$ is the graph of a Hermitian matrix if and only if $\mathcal{L}$ is transversal to ${\mathcal{G}_{\infty}}$.
The Duistermaat index [9, 20] is an integer symplectic invariant of a triple of Lagrangian planes, describing their relative position. Intuitively, it can be thought of as the maximal dimension of a subspace of $\mathcal{L}_{3}$ that lies between $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$. The formal definition we propose differs from that of [9] by taking an axiomatic approach.
Theorem 1.1.
There exists a unique function $\operatorname{\iota}:\Lambda(n)^{3}\to\mathbb{Z}$, called the Duistermaat index, that satisfies the following properties.
  1. (1)
    (Normalization) For any Hermitian matrix $A$,
    $\operatorname{\iota}({\mathcal{G}_{0}},\mathcal{G}_{A},{\mathcal{G}_{\infty}})=n_{-}(A).$ (1.8)
  2. (2)
    (Symplectic Invariance) For any symplectic transformation $\mathsf{S}$ on $\mathbb{C}^{n}\oplus\mathbb{C}^{n}$,
    $\operatorname{\iota}(\mathsf{S}\mathcal{L}_{1},\mathsf{S}\mathcal{L}_{2},\mathsf{S}\mathcal{L}_{3})=\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3}).$ (1.9)
  3. (3)
    (Cocycle Property) For any Lagrangian planes $\mathcal{L}_{1}$, $\mathcal{L}_{2}$, $\mathcal{L}_{3}$ and $\mathcal{L}_{4}$,
    $\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})-\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{4})+\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{3},\mathcal{L}_{4})-\operatorname{\iota}(\mathcal{L}_{2},\mathcal{L}_{3},\mathcal{L}_{4})=0.$ (1.10)
The above index takes values between 00 and $n$:
$0\leq\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})\leq n\qquad\text{for any }\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3}.$ (1.11)
Remark 1.2.
To motivate the terminology in (3), we recall the coboundary operator $\delta$ (see [11, Chapter 3]) which maps functions of $k-1$ arguments to functions of $k$ argument by
$\big{(}\delta\phi\big{)}(x_{1},\ldots,x_{k})=\sum_{j=1}^{k}(-1)^{k-j}\phi(x_{1},\ldots,\widehat{x}_{j},\ldots,x_{k}),$ (1.12)
where a hat indicates the argument has been omitted. Then (1.10) says that the coboundary of $\operatorname{\iota}$ is zero, i.e., $\operatorname{\iota}$ is a cocycle.
Remark 1.3.
Theorem 1.1 will be established in Section 2 by supplying an explicit and self-contained, if somewhat cumbersome, formula for the index. However, the axioms already provide us with a computational algorithm for the index, which also establishes its uniqueness. First we reduce the computation to the case when $\mathcal{L}_{3}$ is transversal to both $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ by using the cocycle identity with some $\mathcal{L}_{4}$ that is transversal to the other three. We can then find a symplectic transformation that maps $\mathcal{L}_{3}$ to ${\mathcal{G}_{\infty}}$ and $\mathcal{L}_{1}$ to ${\mathcal{G}_{0}}$. Since this transformation preserves transversality, the transformed $\mathcal{L}_{2}$ can be represented as the graph of a Hermitian matrix, and (1.8) can be used.
To give an illustration of Theorem 1.1, we prove a useful formula that generalizes (1.8) and also proves (1.2) by setting $A=0$ and relabelling $C$ as $A$.
Theorem 1.4.
For any $n\times n$ Hermitian matrices $A$, $B$ and $C$,
$\operatorname{\iota}(\mathcal{G}_{A},\mathcal{G}_{B},\mathcal{G}_{C})$ $=n_{-}(B-A)-n_{-}(C-A)+n_{-}(C-B)$ (1.13)
$=\left(\delta\widetilde{n}_{-}\right)(A,B,C),$
where $\widetilde{n}_{-}(X,Y):=n_{-}(Y-X)$.
The plan for the rest of the note is as follows: in Section 2 we show that an index satisfying the conditions in Theorem 1.1 exists and in Section 3 we establish some further properties of this index. In Section 4 we interpret the Duistermaat index as the minimal value of the Maslov index [^2] Our definition of the Maslov index, which is reviewed in Section 4, follows the normalization convention of [20]; see [4, Sec. A.4] for a comparison with other conventions in the literature. over non-decreasing paths, namely
$\operatorname{\iota}(\mathcal{L}_{0},\mathcal{L}_{1},\mathcal{M})=\min\left\{\operatorname{Mas}_{[0,1]}\!\big{(}\mathcal{L}(t),\mathcal{M}\big{)}\Big{|}\ \begin{matrix}\mathcal{L}(0)=\mathcal{L}_{0},\mathcal{L}(1)=\mathcal{L}_{1},\\ t\mapsto\mathcal{L}(t)\text{\ is non-decreasing}\end{matrix}\right\}.$ (1.14)
Finally, in Section 5, we prove Theorem 1.4, which we use to deduce (1.1) and its extensions from equation (1.2). We also give an elementary derivation of (1.1) using the Haynsworth inertia formula.
Acknowledgements
This material is based upon work (G.B.) supported by a grant from the Institute for Advanced Study (IAS), School of Mathematics. G.B. also acknowledges support from the NSF Grant DMS-2247473 and from the Association of Former Students of Texas A&M University (through the Faculty Development Leave program that supported the leave to IAS). G.C. acknowledges the support of NSERC grant RGPIN-2017-04259. Y.L. was supported by the NSF grant DMS-2108983/2106157, and would like to thank the Courant Institute of Mathematical Sciences at NYU and especially Prof. Lai-Sang Young for their hospitality. S.S. was supported in part by the NSF Grant DMS-2418900, by the Simons Foundation grant MP-TSM-00002897, and by the Office of the Vice President for Research & Economic Development (OVPRED) at Auburn University through the Research Support Program Grant.
2. Existence of the Duistermaat index
In Remark 1.3 we observed that there is at most one function $\operatorname{\iota}$ satisfying the conditions of Theorem 1.1. The following result establishes existence of such an $\operatorname{\iota}$.
Theorem 2.1.
Given a triple $(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})$ of Lagrangian planes, define a Hermitian form $\Omega$ on $\mathcal{L}_{1}\oplus\mathcal{L}_{2}\oplus\mathcal{L}_{3}$ by
$\Omega(x_{1},x_{2},x_{3};y_{1},y_{2},y_{3})$ $=\omega(x_{1},y_{2})+\omega(x_{2},y_{3})+\omega(x_{3},y_{1})$ (2.1)
$\qquad-\omega(x_{2},y_{1})-\omega(x_{3},y_{2})-\omega(x_{1},y_{3}).$
Then the index $\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})$ defined by
$\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3}):=n_{-}(\Omega)-n+\dim\mathcal{L}_{1}\cap\mathcal{L}_{3}$ (2.2)
satisfies the conditions of Theorem 1.1.
Remark 2.2.
This is not the same as the definition of $\operatorname{\iota}$ given in [9, 20], which requires choosing an arbitrary fourth Lagrangian plane. However, since that definition satisfies the axioms given in Theorem 1.1, it must agree with the index defined by (2.2). One benefit of equation (2.2) is that $\operatorname{\iota}$ is evidently a symplectic invariant. This definition is not optimal for practical computations, however, since $\Omega$ is defined on an $3n$-dimensional space. One also notices a lack of symmetry caused by the $\dim\mathcal{L}_{1}\cap\mathcal{L}_{3}$ term. Both of these concerns are addressed by a more practical (and more symmetric) formula given in Proposition 2.3 below.
Before proving the theorem, we recall that any $n$-dimensional subspace $\mathcal{M}\subset\mathbb{C}^{n}\oplus\mathbb{C}^{n}$ can be described by a frame, which is an injective linear operator
$Z={\begin{pmatrix}X\\ Y\end{pmatrix}}\colon\mathbb{C}^{n}\to\mathbb{C}^{n}\oplus\mathbb{C}^{n}$ (2.3)
whose range is $\mathcal{M}$. Moreover, $\mathcal{M}$ is Lagrangian if and only if $X^{*}Y=Y^{*}X$.
This description is not unique, but it is easy to see that frames $Z$ and $\tilde{Z}$ describe the same subspace if and only if $\tilde{Z}=ZC$ for some invertible $C\colon\mathbb{C}^{n}\to\mathbb{C}^{n}$. Therefore, the set
$E(\mathcal{M}):=\{\epsilon\in\mathbb{R}:X+\epsilon Y\text{ is not invertible}\}$ (2.4)
and the so-called $\epsilon$-Robin map $R^{\epsilon}:\mathbb{C}^{n}\to\mathbb{C}^{n}$,
$\𝐸R^{\epsilon}:=Y(X+\epsilon Y)^{-1},\qquad\epsilon\in\mathbb{R}\backslash E(\mathcal{M}),$ (2.5)
are independent of the choice of frame. The set $E(\mathcal{M})$ is finite, since $\det(X+\epsilon Y)$ is a polynomial in $\epsilon$ that is not identically zero. [^3] Setting $\epsilon=i$ we have $\ker(X+iY)=\ker(X+iY)^{*}(X+iY)=\ker(X^{*}X+Y^{*}Y)=\ker X\cap\ker Y=\{0\}$ because $Z$ is injective. When $\mathcal{M}$ is Lagrangian, the condition $X^{*}Y=Y^{*}X$ implies that
$(X+\epsilon Y)^{*}R^{\epsilon}(X+\epsilon Y)=X^{*}Y+\epsilon Y^{*}Y$ (2.6)
is Hermitian, therefore $R^{\epsilon}$ is Hermitian for all $\𝐸\epsilon\in\mathbb{R}\backslash E(\mathcal{M})$.
Proposition 2.3.
Let $\mathcal{L}_{1}$, $\mathcal{L}_{2}$ and $\mathcal{L}_{3}$ be Lagrangian planes and $R^{\epsilon}_{1}$, $R^{\epsilon}_{2}$ and $R^{\epsilon}_{3}$ the corresponding $\epsilon$-Robin maps, well-defined for any real $\epsilon$ outside the finite set $E_{123}:=E(\mathcal{L}_{1})\cup E(\mathcal{L}_{2})\cup E(\mathcal{L}_{3})$. Then, for the index $\operatorname{\iota}$ defined in (4.3) and any $\epsilon\in\mathbb{R}\backslash E_{123}$, we have
$\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})$ $=n_{-}\Big{(}R_{2}^{\epsilon}-R_{1}^{\epsilon}\Big{)}-n_{-}\Big{(}R_{3}^{\epsilon}-R_{1}^{\epsilon}\Big{)}+n_{-}\Big{(}R_{3}^{\epsilon}-R_{2}^{\epsilon}\Big{)}$ (2.7)
$=\left(\delta\widetilde{n}_{-}\right)\left(R_{1}^{\epsilon},R_{2}^{\epsilon},R_{3}^{\epsilon}\right),$ (2.8)
where $\widetilde{n}_{-}(X,Y):=n_{-}(Y-X)$.
The proof of Theorem 1.1, which we give below, is an easily consequence of this proposition.
Proof.
We follow the proof of a similar identity in [16, Lem. 8.3] (attributed therein to Brian Feldstein). We first compute the form $\Omega$ defined in (2.1) as a Hermitian form on $\mathbb{C}^{n}\oplus\mathbb{C}^{n}\oplus\mathbb{C}^{n}$, where each $\mathbb{C}^{n}$ parametrizes the corresponding Lagrangian plane $\mathcal{L}_{j}$ via a frame $(X_{j},Y_{j})$. If, for instance,
$x_{1}=\begin{pmatrix}X_{1}u_{1}\\ Y_{1}u_{1}\end{pmatrix}\in\mathcal{L}_{1},\qquad y_{2}=\begin{pmatrix}X_{2}v_{2}\\ Y_{2}v_{2}\end{pmatrix}\in\mathcal{L}_{2},$ (2.9)
then
$\omega(x_{1},y_{2})=\left<u_{1},(X_{1}^{*}Y_{2}-Y_{1}^{*}X_{2})v_{2}\right>_{\mathbb{C}^{n}}.$ (2.10)
Similar calculations for the other terms in (2.1) result in $\Omega$ being represented by the Hermitian matrix
$W=\begin{pmatrix}0&X_{1}^{*}Y_{2}-Y_{1}^{*}X_{2}&Y_{1}^{*}X_{3}-X_{1}^{*}Y_{3}\\ Y_{2}^{*}X_{1}-X_{2}^{*}Y_{1}&0&X_{2}^{*}Y_{3}-Y_{2}^{*}X_{3}\\ X_{3}^{*}Y_{1}-Y_{3}^{*}X_{1}&Y_{3}^{*}X_{2}-X_{3}^{*}Y_{2}&0\end{pmatrix}.$ (2.11)
We now note that for $\epsilon\in\mathbb{R}\backslash E_{123}$, and $j,k\in\{1,2,3\}$,
$X_{j}^{*}Y_{k}-Y_{j}^{*}X_{k}$ $=(X_{j}+\epsilon Y_{j})^{*}Y_{k}-Y_{j}^{*}(X_{k}+\epsilon Y_{k})$
$=(X_{j}+\epsilon Y_{j})^{*}(R_{k}^{\epsilon}-R_{j}^{\epsilon})(X_{k}+\epsilon Y_{k}),$ (2.12)
where we used the fact that $R^{\epsilon}$ is Hermitian. We now verify by direct computation that
$W=\frac{1}{2}D^{*}T^{*}\begin{pmatrix}R_{3}^{\epsilon}-R_{2}^{\epsilon}&0&0\\ 0&R_{1}^{\epsilon}-R_{3}^{\epsilon}&0\\ 0&0&R_{2}^{\epsilon}-R_{1}^{\epsilon}\end{pmatrix}TD,$ (2.13)
where the matrices
$T:=\begin{pmatrix}-I&\phantom{-}I&\phantom{-}I\\ \phantom{-}I&-I&\phantom{-}I\\ \phantom{-}I&\phantom{-}I&-I\end{pmatrix},\qquad D:=\begin{pmatrix}X_{1}+\epsilon Y_{1}&0&0\\ 0&X_{2}+\epsilon Y_{2}&0\\ 0&0&X_{3}+\epsilon Y_{3}\end{pmatrix},$ (2.14)
are invertible for $\epsilon\in\mathbb{R}\backslash E_{123}$.
We next observe that, by (2), (2.10), and the definition of a Lagrangian plane in (1.5),
$\ker\left(R_{k}^{\epsilon}-R_{j}^{\epsilon}\right)$ $=\ker\big{(}X_{j}^{*}Y_{k}-Y_{j}^{*}X_{k}\big{)}$
=$=\big{\{}v\in\mathbb{C}^{n}\colon 0=\left<u,(X_{j}^{*}Y_{k}-Y_{j}^{*}X_{k})v\right>_{\mathbb{C}^{n}}\text{ for all }u\in\mathbb{C}^{n}\big{\}}$
$\simeq\big{\{}y\in\mathcal{L}_{k}\colon\omega(x,y)=0\text{ for all }x\in\mathcal{L}_{j}\big{\}}$
$=\mathcal{L}_{j}\cap\mathcal{L}_{k},$ (2.15)
therefore
$n_{0}\left(R_{k}^{\epsilon}-R_{j}^{\epsilon}\right)=\dim\mathcal{L}_{j}\cap\mathcal{L}_{k},$ (2.16)
where $n_{0}(A)=\dim\ker A$.
Finally, applying Sylvester’s law of inertia to (2.13), we use the identity
$n_{-}(R_{1}^{\epsilon}-R_{3}^{\epsilon})=n_{+}(R_{3}^{\epsilon}-R_{1}^{\epsilon})=n-n_{0}(R_{3}^{\epsilon}-R_{1}^{\epsilon})-n_{-}(R_{3}^{\epsilon}-R_{1}^{\epsilon}),$ (2.17)
along with (2.16), to obtain
$n_{-}(\Omega)=n-\dim\mathcal{L}_{1}\cap\mathcal{L}_{3}+n_{-}\left(R_{3}^{\epsilon}-R_{2}^{\epsilon}\right)-n_{-}\left(R_{3}^{\epsilon}-R_{1}^{\epsilon}\right)+n_{-}\left(R_{2}^{\epsilon}-R_{1}^{\epsilon}\right),$ (2.18)
which is equivalent to (2.7). ∎
Remark 2.4.
For future use we observe that (2.13) together with (2.16) gives
$n_{0}(\Omega)=\dim\mathcal{L}_{1}\cap\mathcal{L}_{2}+\dim\mathcal{L}_{1}\cap\mathcal{L}_{3}+\dim\mathcal{L}_{2}\cap\mathcal{L}_{3}.$ (2.19)
Proof of Theorem 2.1.
To check the normalization condition (1.8), we apply (2.7) with $\mathcal{L}_{1}={\mathcal{G}_{0}}$, $\mathcal{L}_{2}=\mathcal{G}_{A}$ and $\mathcal{L}_{3}={\mathcal{G}_{\infty}}$. For small positive $\epsilon$ we have $R_{1}^{\epsilon}=0$, $R_{2}^{\epsilon}=A+O(\epsilon)$ and $R_{3}^{\epsilon}=\epsilon^{-1}I$, therefore
$n_{-}\big{(}R^{\epsilon}_{2}-R^{\epsilon}_{1}\big{)}=n_{-}(A),\qquad n_{-}\big{(}R^{\epsilon}_{3}-R^{\epsilon}_{2}\big{)}=0,\qquad n_{-}\big{(}R^{\epsilon}_{3}-R^{\epsilon}_{1}\big{)}=0.$ (2.20)
The normalization condition then follows from (2.7).
Symplectic invariance of $\operatorname{\iota}$ follows from the symplectic invariance of every term in (2.2), and the cocycle property $\delta\!\operatorname{\iota}=0$ follows from (2.8) by the well-known identity $\delta^{2}\equiv 0$ for the coboundary operator, cf. Remark 1.2.
Finally, to estimate the range of the possible values of the index, we use a symplectic transformation $\mathsf{S}$ such that $\mathsf{S}\mathcal{L}_{3}=\mathcal{G}_{\infty}$ to reduce the question to bounding the range of $\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{G}_{\infty})$. That the latter is between 00 and $n$ follows from a version of equation (2.7):
$\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{G}_{\infty})=n_{-}\Big{(}R_{2}^{\epsilon}-R_{1}^{\epsilon}\Big{)},\qquad\text{for all }0<\epsilon\ll 1.$ (2.21)
Equation (2.21) was already established in [4, Eq. (3.32)], but we repeat the proof here for completeness. We use equation (2.7) with $R^{\epsilon}_{3}=\epsilon^{-1}I$, and the desired result follows once we establish that $n_{-}\left(R_{3}^{\epsilon}-R_{1}^{\epsilon}\right)=0=n_{-}\left(R_{3}^{\epsilon}-R_{2}^{\epsilon}\right)$. In other words, we claim that for any Lagrangian frame $(X,Y)^{T}$, the Hermitian operator $\epsilon^{-1}I-Y(X+\epsilon Y)^{-1}$ is non-negative definite for small enough $\epsilon>0$. It is equivalent to consider
$\epsilon(X+\epsilon Y)^{*}\left(\epsilon^{-1}I-Y(X+\epsilon Y)^{-1}\right)(X+\epsilon Y)=X^{*}X+\epsilon Y^{*}X.$ (2.22)
The right-hand side is identically zero on $\ker(X^{*}X)=\ker X$, while on $\ker(X^{*}X)^{\perp}$ it is a Hermitian perturbation of a positive definite matrix. ∎
3. Further properties of the Duistermaat index
We now derive useful properties of the Duistermaat index, some of which involve algebraic operations that can be defined when Lagrangian subspaces are viewed as self-adjoint linear relations.
3.1. General identities
Theorem 3.1.
The Duistermaat index has the following properties.
  1. (1)
    (Special values) For any Lagrangian $\mathcal{L}$ and $\mathcal{M}$,
    $\operatorname{\iota}(\mathcal{L},\mathcal{L},\mathcal{M})=\operatorname{\iota}(\mathcal{M},\mathcal{L},\mathcal{L})=0,\qquad\operatorname{\iota}(\mathcal{L},\mathcal{M},\mathcal{L})=n-\dim\mathcal{L}\cap\mathcal{M}.$ (3.1)
  2. (2)
    (Permuting the arguments)
    $\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})+\operatorname{\iota}(\mathcal{L}_{2},\mathcal{L}_{1},\mathcal{L}_{3})$ $=n-\dim\mathcal{L}_{1}\cap\mathcal{L}_{2}$ (3.2)
    $\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})+\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{3},\mathcal{L}_{2})$ $=n-\dim\mathcal{L}_{2}\cap\mathcal{L}_{3},$ (3.3)
    $\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})+\operatorname{\iota}(\mathcal{L}_{3},\mathcal{L}_{2},\mathcal{L}_{1})$ $=n-\dim\mathcal{L}_{1}\cap\mathcal{L}_{2}-\dim\mathcal{L}_{2}\cap\mathcal{L}_{3}+\dim\mathcal{L}_{1}\cap\mathcal{L}_{3},$ (3.4)
    $\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})-\dim\mathcal{L}_{1}\cap\mathcal{L}_{3}$ $=\operatorname{\iota}(\mathcal{L}_{3},\mathcal{L}_{1},\mathcal{L}_{2})-\dim\mathcal{L}_{2}\cap\mathcal{L}_{3}$ (3.5)
    $=\operatorname{\iota}(\mathcal{L}_{2},\mathcal{L}_{3},\mathcal{L}_{1})-\dim\mathcal{L}_{1}\cap\mathcal{L}_{2}$ (3.6)
  3. (3)
    (Symplectic additivity) For any three Lagrangians $\mathcal{L}_{1}$, $\mathcal{L}_{2}$, $\mathcal{L}_{3}$ in a symplectic space $V$ and any three Lagrangians $\mathcal{M}_{1}$, $\mathcal{M}_{2}$, $\mathcal{M}_{3}$ in a symplectic space $W$,
    $\operatorname{\iota}(\mathcal{L}_{1}\oplus\mathcal{M}_{1},\mathcal{L}_{2}\oplus\mathcal{M}_{2},\mathcal{L}_{3}\oplus\mathcal{M}_{3})=\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})+\operatorname{\iota}(\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3}).$ (3.7)
  4. (4)
    (Anti-symplectic transformation) If $\mathsf{A}$ is an anti-symplectic linear transformation on $(\mathbb{C}^{n}\oplus\mathbb{C}^{n},\omega)$, i.e., a bijection that satisfies $\omega(\mathsf{A}x,\mathsf{A}y)=-\omega(x,y)$, then
    $\operatorname{\iota}(\mathsf{A}\mathcal{L}_{1},\mathsf{A}\mathcal{L}_{2},\mathsf{A}\mathcal{L}_{3})=\operatorname{\iota}(\mathcal{L}_{3},\mathcal{L}_{2},\mathcal{L}_{1}).$ (3.8)
Proof.
The first two equalities in (3.1) follow directly from (2.7); for the third we also use (2.17) and (2).
To obtain (3.2), we apply the cocycle property
00 0 $=(\delta\!\operatorname{\iota})(\mathcal{L}_{2},\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})$ (3.9)
$=\operatorname{\iota}(\mathcal{L}_{2},\mathcal{L}_{1},\mathcal{L}_{2})-\operatorname{\iota}(\mathcal{L}_{2},\mathcal{L}_{1},\mathcal{L}_{3})+\operatorname{\iota}(\mathcal{L}_{2},\mathcal{L}_{2},\mathcal{L}_{3})-\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})$
and use $\operatorname{\iota}(\mathcal{L}_{2},\mathcal{L}_{1},\mathcal{L}_{2})=n-\dim\mathcal{L}_{1}\cap\mathcal{L}_{2}$ and $\operatorname{\iota}(\mathcal{L}_{2},\mathcal{L}_{2},\mathcal{L}_{3})=0$. Equation (3.3) follows by a similar computation of $(\delta\!\operatorname{\iota})(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3},\mathcal{L}_{2})$; other permutations are obtained by composition of the established two.
Symplectic additivity (3.7) follows by noting that a possible choice of frame for $\mathcal{L}\oplus\mathcal{M}$ is the direct sum of frames. Therefore, for all but finitely many $\epsilon$, one has $R^{\epsilon}_{\mathcal{L}\oplus\mathcal{M}}=R^{\epsilon}_{\mathcal{L}}\oplus R^{\epsilon}_{\mathcal{M}}$. Equation (3.7) then follows from (2.7) and additivity of $n_{-}$ under direct sums.
Finally, to obtain (3.8), we note that the form $\Omega$ of (2.1) changes sign under an anti-symplectic transformation. Therefore, by (2.2),
$\operatorname{\iota}(\mathsf{A}\mathcal{L}_{1},\mathsf{A}\mathcal{L}_{2},\mathsf{A}\mathcal{L}_{3})=n_{-}(-\Omega)-n+\dim\mathcal{L}_{1}\cap\mathcal{L}_{3}.$ (3.10)
We now use $n_{-}(-\Omega)=3n-n_{0}(\Omega)-n_{-}(\Omega)$, as well as (2.19) and (3.4), to get the result. ∎
3.2. Self-adjoint linear relations
For a given splitting of the symplectic space (in our case, $\mathbb{C}^{n}\oplus\mathbb{C}^{n}$), every Lagrangian plane can be viewed as a self-adjoint linear relation; see [18, Sec. 14.1] or [7, Sec. 4.2]. In this context, the difference of two Lagrangian planes, $\mathcal{L}$ and $\mathcal{M}$, is the Lagrangian plane
:=$\mathcal{L}-\mathcal{M}:=\left\{\begin{pmatrix}u\\ v_{{}_{\mathcal{L}}}-v_{{}_{\mathcal{M}}}\end{pmatrix}\in\mathbb{C}^{n}\oplus\mathbb{C}^{n}:\begin{pmatrix}u\\ v_{{}_{\mathcal{L}}}\end{pmatrix}\in\mathcal{L},\ \begin{pmatrix}u\\ v_{{}_{\mathcal{M}}}\end{pmatrix}\in\mathcal{M}\right\},$ (3.11)
and the inverse of a Lagrangian plane $\mathcal{L}$ is
$\mathcal{L}^{-1}:=\left\{\begin{pmatrix}v\\ u\end{pmatrix}\in\mathbb{C}^{n}\oplus\mathbb{C}^{n}:\begin{pmatrix}u\\ v\end{pmatrix}\in\mathcal{L}\right\}.$ (3.12)
If we define
$\operatorname{dom}\mathcal{L}$ :=$:=\left\{u\in\mathbb{C}^{n}:\begin{pmatrix}u\\ v\end{pmatrix}\in\mathcal{L}\text{ for some }v\in\mathbb{C}^{n}\right\},$ (3.13)
$\operatorname{mul}\mathcal{L}$ :=$:=\left\{v\in\mathbb{C}^{n}:\begin{pmatrix}0\\ v\end{pmatrix}\in\mathcal{L}\right\},$ (3.14)
we have $\operatorname{mul}\mathcal{L}=\left(\operatorname{dom}\mathcal{L}\right)^{\perp}$ and $\mathcal{L}$ can be represented as
=$\mathcal{L}=\big{\{}(x,Lx+y)\colon x\in\operatorname{dom}\mathcal{L},\ y\in\operatorname{mul}\mathcal{L}\big{\}}$ (3.15)
for some Hermitian matrix $L\colon\operatorname{dom}\mathcal{L}\to\operatorname{dom}\mathcal{L}$, see, e.g. [2, Thm. 1.5.1] or [18, Prop. 14.2].
Corollary 3.2.
The Duistermaat index enjoys the following further properties.
  1. (1)
    For any Hermitian matrix $A$,
    $\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})=\operatorname{\iota}(\mathcal{L}_{1}-\mathcal{G}_{A},\mathcal{L}_{2}-\mathcal{G}_{A},\mathcal{L}_{3}-\mathcal{G}_{A}).$ (3.16)
  2. (2)
    For any Lagrangian planes $\mathcal{L}_{1}$, $\mathcal{L}_{2}$ and $\mathcal{L}_{3}$,
    $\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})=\operatorname{\iota}(\mathcal{L}_{3}^{-1},\mathcal{L}_{2}^{-1},\mathcal{L}_{1}^{-1}).$ (3.17)
  3. (3)
    For any Hermitian $A$ and Lagrangian $\mathcal{L}$, represented as in (3.15),
    $\operatorname{\iota}(\mathcal{G}_{A},\mathcal{L},{\mathcal{G}_{\infty}})=n_{-}\big{(}L-A_{\operatorname{dom}\mathcal{L}}\big{)},$ (3.18)
    $\operatorname{\iota}(\mathcal{L},\mathcal{G}_{A},{\mathcal{G}_{\infty}})=n_{-}\big{(}A_{\operatorname{dom}\mathcal{L}}-L\big{)}+\dim\mathcal{L}\cap\mathcal{G}_{\infty},$ (3.19)
    where $A_{\operatorname{dom}\mathcal{L}}$ denotes the compression of $A$ to $\operatorname{dom}\mathcal{L}$, namely
    $A_{\operatorname{dom}\mathcal{L}}:=PA|_{\operatorname{dom}\mathcal{L}},$ (3.20)
    where $P$ is the orthogonal projector in $\mathbb{C}^{n}$ onto $\operatorname{dom}\mathcal{L}$.
Remark 3.3.
It is tempting to drop the condition that the first argument in (3.18) is a graph and conjecture that $\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},{\mathcal{G}_{\infty}})=\mathfrak{n}_{-}(\mathcal{L}_{2}-\mathcal{L}_{1})$ for any Lagrangian planes $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$, where the difference $\mathcal{L}_{2}-\mathcal{L}_{1}$ is computed using (3.11) and the negative index $\mathfrak{n}_{-}$ of a Lagrangian plane $\mathcal{L}$ could be defined as the Morse index of its operator part $L$, see (3.15). This enticing conjecture is unfortunately false, as evidenced by the extra term in (3.19).
Proof.
Part (1) follows from symplectic invariance with
$\mathsf{S}:=\begin{pmatrix}I&0\\ A&I\end{pmatrix},$ (3.21)
which maps $\mathcal{L}-\mathcal{G}_{A}\mapsto\mathcal{L}$.
For Part (2) we note that the mapping $(x,y)\mapsto(y,x)$ on $\mathbb{C}^{n}\oplus\mathbb{C}^{n}$, used in the definition of $\mathcal{L}^{-1}$, is anti-symplectic.
For Part (3) we observe that
$\mathcal{L}-\mathcal{G}_{A}$ =$=\big{\{}(x,Lx-Ax+y)\colon x\in\operatorname{dom}\mathcal{L},\ y\in\operatorname{mul}\mathcal{L}\big{\}}$ (3.22)
=$=\big{\{}(x,(L-PA)x+y^{\prime})\colon x\in\operatorname{dom}\mathcal{L},\ y^{\prime}\in\operatorname{mul}\mathcal{L}\big{\}}.$
Using (3.16), we reduce equations (3.18) and (3.19) to their $A=0$ versions,
$\operatorname{\iota}({\mathcal{G}_{0}},\mathcal{L},{\mathcal{G}_{\infty}})=n_{-}(L),\qquad\operatorname{\iota}(\mathcal{L},{\mathcal{G}_{0}},{\mathcal{G}_{\infty}})=n_{-}(-L)+\dim\mathcal{L}\cap\mathcal{V}.$ (3.23)
We now partition the symplectic space as
$\mathbb{C}^{n}\oplus\mathbb{C}^{n}=\big{(}\operatorname{dom}\mathcal{L}\oplus\operatorname{dom}\mathcal{L}\big{)}\oplus\big{(}\operatorname{mul}\mathcal{L}\oplus\operatorname{mul}\mathcal{L}\big{)}.$ (3.24)
It is straightforward to check that $\omega$ restricted to each space is symplectic and the sum of the restrictions is the original $\omega$. With respect to this partition (and with some abuse of notation), we have ${\mathcal{G}_{0}}={\mathcal{G}_{0}}\oplus{\mathcal{G}_{0}}$, $\mathcal{L}=\mathcal{G}_{L}\oplus{\mathcal{G}_{\infty}}$ and ${\mathcal{G}_{\infty}}={\mathcal{G}_{\infty}}\oplus{\mathcal{G}_{\infty}}$. We then use symplectic additivity (3.7), as well as the special values (3.1), to compute
$\operatorname{\iota}_{\mathbb{C}^{n}}({\mathcal{G}_{0}},\mathcal{L},{\mathcal{G}_{\infty}})$ $=\operatorname{\iota}_{\operatorname{dom}\mathcal{L}}({\mathcal{G}_{0}},\mathcal{G}_{L},{\mathcal{G}_{\infty}})+\operatorname{\iota}_{\operatorname{mul}\mathcal{L}}({\mathcal{G}_{0}},{\mathcal{G}_{\infty}},{\mathcal{G}_{\infty}})$ (3.25)
$=\operatorname{\iota}_{\operatorname{dom}\mathcal{L}}({\mathcal{G}_{0}},\mathcal{G}_{L},{\mathcal{G}_{\infty}})=n_{-}(L),$
$\operatorname{\iota}_{\mathbb{C}^{n}}(\mathcal{L},{\mathcal{G}_{0}},{\mathcal{G}_{\infty}})$ $=\operatorname{\iota}_{\operatorname{dom}\mathcal{L}}(\mathcal{G}_{L},{\mathcal{G}_{0}},{\mathcal{G}_{\infty}})+\operatorname{\iota}_{\operatorname{mul}\mathcal{L}}({\mathcal{G}_{\infty}},{\mathcal{G}_{0}},{\mathcal{G}_{\infty}})$ (3.26)
$=\operatorname{\iota}_{\operatorname{dom}\mathcal{L}}(\mathcal{G}_{L},{\mathcal{G}_{0}},{\mathcal{G}_{\infty}})+\dim\operatorname{mul}\mathcal{L}=n_{-}(-L)+\dim\mathcal{L}\cap{\mathcal{G}_{\infty}},$
which completes the proof. ∎
4. Relationship with other indices
4.1. The Kashiwara index
Another commonly used index of a triple of Lagrangian planes arose independently in the work of Hörmander, Kashiwara and Wall; see, for instance, [19], [13, Eq. (3.3.15)], [14, Def. 1.5.1], [8, Sec. 8], [10, Sec 4.2]. The Hörmander–Kashiwara–Wall index $s(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})$ is defined [^4] Only the latter reference considered complex symplectic space; the signs in our definition of the form $\Omega$ are different. as the signature of the form $\Omega$ from Theorem 2.1,
$s(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3}):=\operatorname{Sig}(\Omega)=n_{+}(\Omega)-n_{-}(\Omega).$ (4.1)
The advantage in dealing with $\operatorname{Sig}(A)$ is its symmetry property $\operatorname{Sig}(-A)=-\operatorname{Sig}(A)$, in contrast to (2.17). However, our motivation comes, ultimately, from operators that are bounded only from below; in this case $n_{-}$ may be well-defined, while $n_{+}$ will not be.
The two indices are connected through the following formulas.
Lemma 4.1.
For any Lagrangian planes $\mathcal{L}_{1}$, $\mathcal{L}_{2}$ and $\mathcal{L}_{3}$,
$s(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})=\operatorname{\iota}(\mathcal{L}_{2},\mathcal{L}_{1},\mathcal{L}_{3})-\operatorname{\iota}(\mathcal{L}_{3},\mathcal{L}_{1},\mathcal{L}_{2})$ (4.2)
or, equivalently,
$\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})=\frac{1}{2}\Big{(}n-\dim\mathcal{L}_{1}\cap\mathcal{L}_{2}+\dim\mathcal{L}_{1}\cap\mathcal{L}_{3}-\dim\mathcal{L}_{2}\cap\mathcal{L}_{3}-s(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})\Big{)}.$ (4.3)
Proof.
Applying (2.7) to $\operatorname{\iota}(\mathcal{L}_{2},\mathcal{L}_{1},\mathcal{L}_{3})$ and $\operatorname{\iota}(\mathcal{L}_{3},\mathcal{L}_{1},\mathcal{L}_{2})$ and recalling that $n_{-}(A)=n_{+}(-A)$, we get
$\operatorname{\iota}(\mathcal{L}_{2},\mathcal{L}_{1},\mathcal{L}_{3})-\operatorname{\iota}(\mathcal{L}_{3},\mathcal{L}_{1},\mathcal{L}_{2})=\operatorname{Sig}(R_{2}^{\epsilon}-R_{1}^{\epsilon})+\operatorname{Sig}(R_{3}^{\epsilon}-R_{1}^{\epsilon})+\operatorname{Sig}(R_{3}^{\epsilon}-R_{2}^{\epsilon}),$ (4.4)
which is equal to $s(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})$ by (4.1) and (2.13). To obtain (4.3) we use Theorem 3.1(2). ∎
It follows that $\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})$ is the only non-trivial invariant of a triple of Lagrangian planes, in the following sense.
Proposition 4.2.
For any two triples $(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})$ and $(\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3})$ of Lagrangian planes, there is a linear symplectic transformation $\mathsf{S}$ satisfying $\mathsf{S}\mathcal{L}_{j}=\mathcal{M}_{j}$ for all $j\in\{1,2,3\}$ if and only if
$\dim\mathcal{L}_{j}\cap\mathcal{L}_{k}=\dim\mathcal{M}_{j}\cap\mathcal{M}_{k},\quad\text{for all }j,k\in\{1,2,3\},$ (4.5)
$\dim\mathcal{L}_{1}\cap\mathcal{L}_{2}\cap\mathcal{L}_{3}=\dim\mathcal{M}_{1}\cap\mathcal{M}_{2}\cap\mathcal{M}_{3},$ (4.6)
$\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})=\operatorname{\iota}(\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3}).$ (4.7)
Proof.
Proposition 4.2 is well-known when (4.7) is replaced by the condition $s(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})=s(\mathcal{M}_{1},\mathcal{M}_{2},\mathcal{M}_{3})$, see, e.g. [16, Thm. 8.4(d)] or [1, Prop. 4.4]. But the indices $s$ and $\operatorname{\iota}$ are related by an invertible linear transformation involving the invariants appearing in (4.5). ∎
Corollary 4.3.
For any two pairs $(\mathcal{L}_{1},\mathcal{L}_{2})$ and $(\mathcal{M}_{1},\mathcal{M}_{2})$ of Lagrangian planes, there is a linear symplectic transformation $\mathsf{S}$ satisfying $\mathsf{S}\mathcal{L}_{j}=\mathcal{M}_{j}$, $j=1,2$, if and only if
$\dim\mathcal{L}_{1}\cap\mathcal{L}_{2}=\dim\mathcal{M}_{1}\cap\mathcal{M}_{2}.$ (4.8)
Proof.
While usually established before the more complicated case of a triple, see [16, Exercise 8.1], this fact also follows from Proposition 4.2 by applying it to the triples $(\mathcal{L}_{1},\mathcal{L}_{1},\mathcal{L}_{2})$ and $(\mathcal{M}_{1},\mathcal{M}_{1},\mathcal{M}_{2})$ and using (3.1). ∎
4.2. The Maslov index
Suppose $\mathcal{L}\colon[0,1]\to\Lambda(n)$ is a differentiable path of Lagrangian planes, given by a frame $\big{(}X(t),Y(t)\big{)}$. For $t_{0}\in[0,1]$ and $v\in\mathcal{L}(t_{0})$ we define the [^5] The equivalence of (4.9) to other common formulas in the literature is proved in [4, Thm. 2.1]. crossing form $\mathfrak{q}$ on $\mathcal{L}(t_{0})$ by
$\mathfrak{q}(v,v)=\left<x,\big{(}X^{*}(t_{0})Y^{\prime}(t_{0})-Y^{*}(t_{0})X^{\prime}(t_{0})\big{)}x\right>_{\mathbb{C}^{n}},\qquad\text{where }v=\begin{pmatrix}X(t_{0})x\\ Y(t_{0})x\end{pmatrix}.$ (4.9)
We say that the path $\mathcal{L}(\cdot)$ is non-decreasing if $\mathfrak{q}$ is positive semidefinite at all $t$. We say that it is regular, with respect to a given Lagrangian plane $\mathcal{M}$, if it is continuously differentable and the restricted quadratic form $\mathfrak{q}|_{\mathcal{L}(t_{0})\cap\mathcal{M}}$ is non-degenerate whenever $\mathcal{L}(t_{0})\cap\mathcal{M}$ is nontrivial. Such $t_{0}$ are referred to as crossings or conjugate points.
For a regular path $\mathcal{L}(\cdot)$ the conjugate points are isolated, and its Maslov index with respect to a Lagrangian reference plane $\mathcal{M}$ is given by
$\operatorname{Mas}_{[0,1]}\!\big{(}\mathcal{L}(t),\mathcal{M}\big{)}:=\sum_{t\in[0,1)}n_{+}(\mathfrak{q}|_{\mathcal{L}(t)\cap\mathcal{M}})-\sum_{t\in(0,1]}n_{-}(\mathfrak{q}|_{\mathcal{L}(t)\cap\mathcal{M}}).$ (4.10)
Remark 4.4.
There are two ways of viewing the formula (4.10), depending on how one defines the Maslov index. One approach is to define the Maslov index for any continuous path via the spectral flow of a certain family of unitary operators. For a regular path one then recovers (4.10) as a special case; see [6, 10] for details. Alternately, one can take (4.10) as the definition of the Maslov index for a regular path, then define the Maslov index of a continuous path by first deforming it to a regular path and then arguing that the resulting index is independent of the choice of deformation, as in [17].
Theorem 4.5.
For arbitrary Lagrangian planes $\mathcal{L}_{0}$, $\mathcal{L}_{1}$ and $\mathcal{M}$ one has
$\operatorname{\iota}(\mathcal{L}_{0},\mathcal{L}_{1},\mathcal{M})=\min\left\{\operatorname{Mas}_{[0,1]}\!\big{(}\mathcal{L}(t),\mathcal{M}\big{)}\Big{|}\ \begin{matrix}\mathcal{L}(0)=\mathcal{L}_{0},\mathcal{L}(1)=\mathcal{L}_{1},\\ t\mapsto\mathcal{L}(t)\text{\ is non-decreasing}\end{matrix}\right\}.$ (4.11)
This theorem highlights the usefulness of the Duistermaat index, since computation of the Maslov index is far more difficult than applying an algebraic identity such as (1.13).
Lemma 4.6.
For any Lagrangian planes $\mathcal{L}_{0}$ and $\mathcal{L}_{1}$, there exists a smooth, non-decreasing path $\mathcal{L}\colon[0,1]\to\Lambda(n)$ between $\mathcal{L}_{0}$ and $\mathcal{L}_{1}$ such that
$\operatorname{\iota}(\mathcal{L}_{0},\mathcal{L}_{1},\mathcal{M})=\operatorname{Mas}_{[0,1]}\!\big{(}\mathcal{L}(t),\mathcal{M}\big{)}$ (4.12)
for any Lagrangian plane $\mathcal{M}$.
Proof.
Let $Q\colon\mathbb{C}^{n}\to\mathbb{C}^{n}$ be an arbitrary orthogonal projector with $\dim\ker Q=\dim\mathcal{L}_{0}\cap\mathcal{L}_{1}$. Since $\mathcal{G}_{0}\cap\mathcal{G}_{Q}$ and $\mathcal{L}_{0}\cap\mathcal{L}_{1}$ have the same dimension, by Corollary 4.3 there exists a symplectic transformation $\mathsf{S}$ so that $\mathsf{S}\mathcal{L}_{0}=\mathcal{G}_{0}$ and $\mathsf{S}\mathcal{L}_{1}=\mathcal{G}_{Q}$. Since both quantities in (4.12) are symplectic invariants, without loss of generality we can assume that $\mathsf{S}$ is identity. That is, it suffices to find a non-decreasing path $\mathcal{L}(\cdot)$ with $\mathcal{L}(0)=\mathcal{G}_{0}$, $\mathcal{L}(1)=\mathcal{G}_{Q}$ and $\operatorname{\iota}(\mathcal{G}_{0},\mathcal{G}_{Q},\mathcal{M})=\operatorname{Mas}\!\big{(}\mathcal{L}(t),\mathcal{M}\big{)}$ for any Lagrangian plane $\mathcal{M}$. We claim that $\mathcal{L}(t)=\mathcal{G}_{tQ}$ has this property.
Representing $\mathcal{M}$ using a Hermitian matrix $M$ on $\operatorname{dom}\mathcal{M}$, as in (3.15), and letting $P$ denote the orthogonal projection onto $\operatorname{dom}\mathcal{M}$, we use the cocycle property (1.10) to obtain
$\operatorname{\iota}(\mathcal{G}_{0},\mathcal{G}_{Q},\mathcal{M})$ $=\operatorname{\iota}(\mathcal{G}_{0},\mathcal{G}_{Q},\mathcal{G}_{\infty})-\operatorname{\iota}(\mathcal{G}_{0},\mathcal{M},\mathcal{G}_{\infty})+\operatorname{\iota}(\mathcal{G}_{Q},\mathcal{M},\mathcal{G}_{\infty})$ (4.13)
$=n_{-}(Q)-n_{-}(M)+n_{-}\big{(}M-Q_{\operatorname{dom}\mathcal{M}}\big{)},$ (4.14)
where we used (1.8) for the first term on the right-hand side and (3.18) for the remaining terms. Since $n_{-}(Q)=0$, the proof of (4.12) will be complete once we show that
$\operatorname{Mas}_{[0,1]}\!\big{(}\mathcal{L}(t),\mathcal{M}\big{)}=n_{-}\big{(}M-Q_{\operatorname{dom}\mathcal{M}}\big{)}-n_{-}(M).$ (4.15)
To calculate this Maslov index we define the two-parameter family $\mathcal{L}(s,t)=\mathcal{G}_{tQ-sI}$. It is clear that $\mathcal{L}(t)=\mathcal{L}(0,t)$ is homotopic, with fixed endpoints, to the concatenation of $\mathcal{L}(s,0)|_{s\in[0,s_{*}]}$, $\mathcal{L}(s_{*},t)|_{t\in[0,1]}$ and $\mathcal{L}(s_{*}-s,1)|_{s\in[0,s_{*}]}$, so the index of $\mathcal{L}(t)$ is the sum of these three indices. We choose the parameter $s_{*}>0$ to be small enough to guarantee that neither $M$ nor $M-Q_{\operatorname{dom}\mathcal{M}}$ has any spectrum in the interval $[-s_{*},0)$.
Recalling the representation of $\mathcal{M}$ in (3.15) and computing as in (3.22), we observe that
$\mathcal{L}(s,t)\cap\mathcal{M}=\big{\{}(x,Mx):x\in\operatorname{dom}\mathcal{L}\text{ and }Mx=tPQx-sx\big{\}}.$ (4.16)
Our condition on $s_{*}$ means that this intersection is trivial for $t=0$ and $s\in(0,s_{*}]$, as well as for $t=1$ and $s\in(0,s_{*}]$.
The path $\mathcal{L}(s,0)|_{[0,s_{*}]}$ may still have a conjugate point at $s=0$. However, the $s$-crossing form, computed via (4.9), is $x,x\right>_{\mathbb{C}^{n}}$, which is negative definite and hence generates no contribution to (4.10) at the initial point of a path. We conclude that $\operatorname{Mas}_{[0,s_{*}]}\!\big{(}\mathcal{L}(s,0),\mathcal{M}\big{)}=0$ and, by similar reasoning, $\operatorname{Mas}_{[0,s_{*}]}\!\big{(}\mathcal{L}(s_{*}-s,1),\mathcal{M}\big{)}=0$.
We now evaluate the remaining term, $\operatorname{Mas}_{[0,1]}\!\big{(}\mathcal{L}(s_{*},t),\mathcal{M}\big{)}$. The $t$-crossing form evaluates to $x,Qx\right>_{\mathbb{C}^{n}}$, which we claim to be strictly positive on $\mathcal{L}(s_{*},t)\cap\mathcal{M}$ whenever this intersection is non-trivial. Indeed, the form is obviously non-negative and if $Qx=0$ for some nonzero $x\in\operatorname{dom}\mathcal{L}$, then (4.16) implies $Mx=-s_{*}x$, which is excluded by our condition that $M$ has no spectrum in $[-s_{*},0)$.
It follows that each $n_{+}$ term in (4.10) is simply the dimension of the space $\mathcal{L}(s_{*},t)\cap\mathcal{M}$,
$\operatorname{Mas}_{[0,1]}\!\big{(}\mathcal{L}(s_{*},t),\mathcal{M}\big{)}=\sum_{t\in[0,1)}\dim\ker\big{(}M-tQ_{\operatorname{dom}\mathcal{M}}+s_{*}I\big{)}.$ (4.17)
Since the eigenvalues of $M-tQ_{\operatorname{dom}\mathcal{M}}+s_{*}I$ are decreasing in $t$, the sum in (4.17) is equal to the number of eigenvalues of $M-tQ_{\operatorname{dom}\mathcal{M}}+s_{*}I$ that cross 0 to become negative, thus
$\operatorname{Mas}_{[0,1]}\!\big{(}\mathcal{L}(s_{*},t),\mathcal{M}\big{)}=n_{-}\big{(}M-Q_{\operatorname{dom}\mathcal{M}}+s_{*}I\big{)}-n_{-}(M+s_{*}I).$ (4.18)
Finally, we observe that our conditions on the smallness of $s_{*}$ imply
$n_{-}\big{(}M-Q_{\operatorname{dom}\mathcal{M}}+s_{*}I\big{)}=n_{-}\big{(}M-Q_{\operatorname{dom}\mathcal{M}}\big{)},\qquad n_{-}(M+s_{*}I)=n_{-}(M),$ (4.19)
establishing (4.15). ∎
Remark 4.7.
If $\ker M\cap\ker Q$ is nontrivial, then this subspace is contained in $\mathcal{L}(0,t)\cap\mathcal{M}$ for all $t$. This means the crossings are not isolated and the crossing form is degenerate, so (4.10) does not apply. This is the reason for the introduction of $s$ in the proof above. A similar homotopy argument was used in [17] to prove the “zero axiom” of the Maslov index, which says that a path $\mathcal{L}(t)$ for which $\dim(\mathcal{L}(t)\cap\mathcal{M})$ is constant dimension has $\operatorname{Mas}\!\big{(}\mathcal{L}(t),\mathcal{M}\big{)}=0$.
As a corollary of Lemma 4.6 we obtain the Zhou–Wu–Zhu identity [20, Thm. 1.1].
Theorem 4.8 (Zhou–Wu–Zhu identity).
For any path $\mathcal{L}\colon[0,1]\to\Lambda(n)$ and any Lagrangian planes $\mathcal{M}_{1}$, $\mathcal{M}_{2}$,
$\operatorname{Mas}_{[0,1]}\!\big{(}\mathcal{L}(t),\mathcal{M}_{1}\big{)}-\operatorname{Mas}_{[0,1]}\!\big{(}\mathcal{L}(t),\mathcal{M}_{2}\big{)}$ $=\operatorname{\iota}\!\big{(}\mathcal{L}(0),\mathcal{L}(1),\mathcal{M}_{1}\big{)}-\operatorname{\iota}\!\big{(}\mathcal{L}(0),\mathcal{L}(1),\mathcal{M}_{2}\big{)}$ (4.20)
$=\operatorname{\iota}\!\big{(}\mathcal{L}(0),\mathcal{M}_{1},\mathcal{M}_{2}\big{)}-\operatorname{\iota}\!\big{(}\mathcal{L}(1),\mathcal{M}_{1},\mathcal{M}_{2}\big{)}.$ (4.21)
Proof.
Use Lemma 4.6 to choose a path ~$\widetilde{\mathcal{L}}\colon[1,2]\to\Lambda(n)$ from $\mathcal{L}(1)$ to $\mathcal{L}(0)$ such that
$\operatorname{Mas}_{[1,2]}\!\big{(}\widetilde{\mathcal{L}}(t),\mathcal{M}_{j}\big{)}=\operatorname{\iota}\!\big{(}\mathcal{L}(1),\mathcal{L}(0),\mathcal{M}_{j}\big{)},\qquad j=1,2.$ (4.22)
The concatenation of $\mathcal{L}$ and $\widetilde{\mathcal{L}}$ is a loop, therefore its Maslov index does not depend on the reference plane $\mathcal{M}_{j}$. Using the additivity of the Maslov index under concatenation, we get
$\operatorname{Mas}_{[0,1]}\!\big{(}\mathcal{L}(t),\mathcal{M}_{1}\big{)}+\operatorname{\iota}\!\big{(}\mathcal{L}(1),\mathcal{L}(0),\mathcal{M}_{1}\big{)}=\operatorname{Mas}_{[0,1]}\!\big{(}\mathcal{L}(t),\mathcal{M}_{2}\big{)}+\operatorname{\iota}\!\big{(}\mathcal{L}(1),\mathcal{L}(0),\mathcal{M}_{2}\big{)}.$ (4.23)
Equation (4.20) now follows by application of (3.2), and equation (4.21) follows from the cocycle property (1.10). ∎
Proof of Theorem 4.5.
For any non-decreasing path $\mathcal{L}(t)$ from $\mathcal{L}_{0}$ to $\mathcal{L}_{1}$, we use (4.20) with $\mathcal{M}_{1}=\mathcal{M}$, $\mathcal{M}_{2}=\mathcal{L}(1)$ and note that $\operatorname{\iota}\!\big{(}\mathcal{L}(0),\mathcal{L}(1),\mathcal{L}(1)\big{)}=0$ by (3.1). We obtain
$\operatorname{\iota}\!\big{(}\mathcal{L}(0),\mathcal{L}(1),\mathcal{M}\big{)}=\operatorname{Mas}_{[0,1]}\!\big{(}\mathcal{L}(t),\mathcal{M}\big{)}-\operatorname{Mas}_{[0,1]}\!\big{(}\mathcal{L}(t),\mathcal{L}(1)\big{)}\leq\operatorname{Mas}_{[0,1]}\!\big{(}\mathcal{L}(t),\mathcal{M}\big{)},$ (4.24)
because the path $\mathcal{L}(\cdot)$ is non-decreasing. Lemma 4.6 concludes the proof. ∎
5. The Morse index of a difference of Hermitian matrices
We start by proving Theorem 1.4, which asserts that
$\operatorname{\iota}(\mathcal{G}_{A},\mathcal{G}_{B},\mathcal{G}_{C})=n_{-}(B-A)-n_{-}(C-A)+n_{-}(C-B)$ (5.1)
for any $n\times n$ Hermitian matrices $A$, $B$ and $C$.
Proof of Theorem 1.4.
We first use the cocycle property to write
$\operatorname{\iota}(\mathcal{G}_{A},\mathcal{G}_{B},\mathcal{G}_{C})=\operatorname{\iota}(\mathcal{G}_{A},\mathcal{G}_{B},{\mathcal{G}_{\infty}})-\operatorname{\iota}(\mathcal{G}_{A},\mathcal{G}_{C},{\mathcal{G}_{\infty}})+\operatorname{\iota}(\mathcal{G}_{B},\mathcal{G}_{C},{\mathcal{G}_{\infty}}).$ (5.2)
Next, we observe that for any Hermitian $A$ and $B$, the symplectic transformation
$\mathsf{S}:=\begin{pmatrix}I&0\\ A&I\end{pmatrix}$ (5.3)
maps ${\mathcal{G}_{0}}\mapsto\mathcal{G}_{A}$, $\mathcal{G}_{B-A}\mapsto\mathcal{G}_{B}$, and ${\mathcal{G}_{\infty}}\mapsto{\mathcal{G}_{\infty}}$. Symplectic invariance and the normalization condition (1.8) together imply
$\operatorname{\iota}(\mathcal{G}_{A},\mathcal{G}_{B},{\mathcal{G}_{\infty}})=\operatorname{\iota}({\mathcal{G}_{0}},\mathcal{G}_{B-A},{\mathcal{G}_{\infty}})=n_{-}(B-A).$ (5.4)
Applying this to all three terms on the right-hand side of (5.2) completes the proof. ∎
We are now ready to prove equation (1.1) and its variants. For a Hermitian matrix $X$, we let $X^{+}$ denote the Moore–Penrose pseudoinverse and let $P_{X}$ denote the orthogonal projector onto $\operatorname{Ran}X$.
Proposition 5.1.
Let $A$ and $B$ be $n\times n$ Hermitian matrices.
  1. (1)
    If $\ker A\subseteq\ker B$, then
    $n_{-}\big{(}A-B\big{)}=n_{-}(A)-n_{-}(B)+n_{-}\big{(}B^{+}-P_{B}A^{+}P_{B}\big{)}.$ (5.5)
  2. (2)
    If $\ker B\subseteq\ker A$, then
    $n_{-}\big{(}A-B\big{)}=n_{-}(A)-n_{-}(B)+n_{-}\big{(}P_{A}B^{+}P_{A}-A^{+}\big{)}+n_{0}(A)-n_{0}(B).$ (5.6)
We note that equation (1.1) follows from either one of (5.5) and (5.6): for an invertible $X$, $P_{X}$ is the identity and $X^{+}=X^{-1}$.
Proof.
In the case when $\ker A\subseteq\ker B$, the Hermitian matrices $A$, $B$, $A-B$ and $B^{+}-P_{B}A^{+}P_{B}$ are all identically zero on $\ker A$. This means we can restrict $A$ and $B$ (and thus $A-B$ and $B^{+}-P_{B}A^{+}P_{B}$) to the subspace $(\ker A)^{\perp}$, on which $A$ is invertible, without changing their Morse indices. It thus suffices to prove (5.5) when $A$ is invertible.
Using Theorem 1.4, we just need to show that $\operatorname{\iota}(\mathcal{G}_{0},\mathcal{G}_{B},\mathcal{G}_{A})=n_{-}\big{(}B^{+}-P_{B}A^{+}P_{B}\big{)}$. We use (3.17) to get
$\operatorname{\iota}(\mathcal{G}_{0},\mathcal{G}_{B},\mathcal{G}_{A})=\operatorname{\iota}(\mathcal{G}_{A}^{-1},\mathcal{G}_{B}^{-1},{\mathcal{G}_{\infty}}),$ (5.7)
and observe that for invertible $A$ we have
$\mathcal{G}_{A}^{-1}=\mathcal{G}_{A^{-1}}=\mathcal{G}_{A^{+}},\qquad\mathcal{G}_{B}^{-1}=\big{\{}(x,B^{+}x+y)\colon x\in\operatorname{Ran}B,\ y\in\ker B\big{\}},$ (5.8)
cf. (3.15). The desired conclusion now follows from equation (3.18).
The proof of (5.6) proceeds similarly, by restricting all matrices in the equation to $(\ker B)^{\perp}$ and observing that
$n_{0}(A)-n_{0}(B)=n_{0}\left(A|_{(\ker B)^{\perp}}\right)-n_{0}\left(B|_{(\ker B)^{\perp}}\right)=n_{0}\left(A|_{(\ker B)^{\perp}}\right).$ (5.9)
To establish (5.6) under the assumption that $B$ is invertible, we use Theorem 1.4, equation (5.7), an analogue of (5.8) with $A$ and $B$ swapped, and, finally, (3.19). ∎
Corollary 5.2.
For any invertible $n\times n$ Hermitian matrices $A$ and $B$,
$n_{-}\big{(}A+B\big{)}+n_{0}\big{(}A+B\big{)}+n_{-}\big{(}A^{-1}+B^{-1}\big{)}=n_{-}(A)+n_{-}(B).$ (5.10)
Proof.
Equation (1.1) with $-A$ in place of $A$ becomes
$n_{+}\big{(}A+B\big{)}-n_{-}\big{(}A^{-1}+B^{-1}\big{)}=n_{+}(A)-n_{-}(B).$ (5.11)
We then use $n_{+}(X)=n-n_{-}(X)-n_{0}(X)$ to obtain (5.10); $n_{0}(A)$ disappears because $A$ is invertible. ∎
Similar computations lead to a much simpler proof of [4, Prop. 6.1], which we give below.
Proposition 5.3 (Prop. 6.1 of [4]).
Assume $\mathcal{L}_{1}$, $\mathcal{L}_{2}$ and $\mathcal{L}_{3}$ are Lagrangian planes such that $\mathcal{L}_{3}$ is transversal to $\mathcal{L}_{1}$, $\mathcal{L}_{2}$ and ${\mathcal{G}_{\infty}}$. Then
$\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})=n_{-}\left((\mathcal{L}_{1}-\mathcal{L}_{3})^{-1}-(\mathcal{L}_{2}-\mathcal{L}_{3})^{-1}\right).$ (5.12)
Proof.
We have the following chain of equalities
$\operatorname{\iota}(\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3})$ $=\operatorname{\iota}(\mathcal{L}_{3},\mathcal{L}_{1},\mathcal{L}_{2})$ (5.13)
$=\operatorname{\iota}({\mathcal{G}_{0}},\mathcal{L}_{1}-\mathcal{L}_{3},\mathcal{L}_{2}-\mathcal{L}_{3})$ (5.14)
$=\operatorname{\iota}\left((\mathcal{L}_{2}-\mathcal{L}_{3})^{-1},(\mathcal{L}_{1}-\mathcal{L}_{3})^{-1},{\mathcal{G}_{\infty}}\right)$ (5.15)
$=n_{-}\left((\mathcal{L}_{1}-\mathcal{L}_{3})^{-1}-(\mathcal{L}_{2}-\mathcal{L}_{3})^{-1}\right).$ (5.16)
In the first step we applied (3.5) and used $\mathcal{L}_{1}\cap\mathcal{L}_{3}=\mathcal{L}_{2}\cap\mathcal{L}_{3}=\{0\}$. In the second step we used (3.16), which is allowed because $\mathcal{L}_{3}$ is transversal to ${\mathcal{G}_{\infty}}$ and therefore is a graph of a matrix. In the third step we used (3.17). We remark that because $\mathcal{L}_{3}$ is transversal to $\mathcal{L}_{1}$, $\mathcal{L}_{1}-\mathcal{L}_{3}$ is transversal to ${\mathcal{G}_{0}}$ and thus $(\mathcal{L}_{1}-\mathcal{L}_{3})^{-1}$ is transversal to ${\mathcal{G}_{\infty}}$. Similarly, $(\mathcal{L}_{2}-\mathcal{L}_{3})^{-1}$ is also a graph. Therefore, in the last step, we applied (3.16) to subtract $(\mathcal{L}_{2}-\mathcal{L}_{3})^{-1}$ and used the normalization condition (1.8). ∎
Remark 5.4.
The identity (1.1) can also be obtained using Haynsworth’s inertia additivity formula. For a block Hermitian matrix $H=\left(\begin{smallmatrix}A&B\\ B^{*}&D\end{smallmatrix}\right)$ with invertible $A$, the Haynsworth formula [12] states
$n_{-}(H)=n_{-}(A)+n_{-}\left(D-B^{*}A^{-1}B\right).$ (5.17)
A similar identity holds if $D$ is invertible. Applying this formula in two possible ways to the matrix $\left(\begin{smallmatrix}A&I\\ I&B^{-1}\end{smallmatrix}\right)$ yields
$n_{-}(A)+n_{-}\left(B^{-1}-A^{-1}\right)$ $=n_{-}\left(B^{-1}\right)+n_{-}\left(A-I\left(B^{-1}\right)^{-1}\right)$ (5.18)
$=n_{-}(B)+n_{-}(A-B),$ (5.19)
which establishes (1.1). Proposition 5.1 can be proved analogously, however, this requires an appropriate extension of the Haynsworth formula such as [15, Thm. 6.1] or [3, Thm. A.1].